# TEXTOC: Text-driven Object-Centric Style Transfer

### v2

@@ 4.1 Implementation Details
We commence our training with an initial set of 20 early iterations and completed a total of 200 iterations using the Adam optimizer [30]. The training began with an initial learning rate of $5\times 10^{-4}$, halved after the first 100 iterations. We incorporate the perspective augmentation function from the PyTorch library [39]. Notably, our model treats each scene independently, processing them on a per-scene basis. On average, the training time for each text prompt is approximately 45 seconds when conducted on an NVIDIA A6000 GPU. For further details on our training procedure, please refer to the supplementary material.
4.2 Evaluation metrics
We conduct a comprehensive quantitative evaluation using a variety of metrics that are well-established in the field of text-driven style transfer as outlined in Tab. 1. The similarity metric (Sim) quantifies the likeness between the image and text embedding vectors, assessing the congruence of the two embeddings. We also apply content metrics (Con) using the VGG content loss [26] and style metrics (Sty) using the mean-variance style loss [22]. These metrics are adapted from traditional calculations but specifically mask either the foreground or background of input images to focus analysis on selected areas. To thoroughly evaluate the quality of the stylized image backgrounds from multiple aspects, we conduct additional assessments specifically targeting the background masks. For this purpose, we utilize the SSIM metric [52], which evaluates structural, luminance, and color attributes. Additionally, we employ the DISTS measure [10] to ensure deep structural and textural consistency in these areas. Furthermore, we use the PSNR [13] metric and absolute difference (L1) for an overall assessment of signal fidelity and basic visual quality in the backgrounds.
[…]
$\begin{gathered}\text{Sim}_{F}=\frac{E_{T}\left(T^{\text{tgt}}_{i}\right)\cdot E_{I}(I^{\text{out}}\odot M^{\text{fg\_gt}})}{|E_{T}\left(T^{\text{tgt}}_{i}\right)|\cdot|E_{I}(I^{\text{out\_f}})|},\\ \text{L1}_{B}=\frac{\sum|I^{\text{src}}\odot M^{\text{bg\_gt}}-I^{\text{out}}\odot M^{\text{bg\_gt}}|}{255\cdot\text{num}(I^{\text{src}}\odot M^{\text{bg\_gt}})},\end{gathered}$ (7)
@@ 4.3 Comparison to state-of-the-arts
In the qualitative evaluation, the first row of Fig. 3 showcases the transformation of a mountain into a volcano using the style text “Volcano”. Our method adeptly preserves the inherent structure of the mountain while seamlessly integrating the volcanic style. In contrast, Text2LIVE often reconstructs an entirely new mountain structure, typically characterized by a singular hole indicative of magma. For the five diffusion-based style transfer models, it is noticeable that the images either do not look realistic or fail to adequately reflect the style of a volcano. This comparison underscores the nuanced capability of our approach in maintaining the original form while applying distinct style transformations. Utilizing source texts like “wooden door”, “croissant” and “red umbrella”, our method demonstrates a precise capture of object locations, contrasting with other models that often indiscriminately apply style transfer to the entire image. While Text2LIVE shows results similar to ours, it notably applies style transfer to unintended areas (e.g., wooden door with “gray marble” as style text) and fail to properly reflect the style indicated by the style text (e.g., red umbrella in ”Starry night by vincent Van Gogh”).
Figure 5: Our stylization results with complex image-text scenarios.
@@ 4.4 Ablation study
To demonstrate the efficacy of our method, we conduct an ablation study, with quantitative results presented in Tab. 2 and qualitative insights in Fig. 4. Notably, the quantitative results in Tab. 2-(e) which incorporates all proposed losses $\mathcal{L}_{\text{dir}}$, $\mathcal{L}_{\text{con}}$, and $\mathcal{L}_{\text{abp}}$, indicate that our method achieves the best performance. Notably, the absence of the adaptive background preservation loss $\mathcal{L}_{\text{abp}}$ in Tab. 2-(c) results in a significant drop in the background similarity metric $\text{Sim}_{B}$. Omitting the patch distribution consistency loss $\mathcal{L}_{\text{con}}$ in Tab. 2-(d) leads to a decrease in performance across all foreground quality metrics. These findings clearly demonstrate the critical role of each proposed loss, affirming their essential contributions to our design.
The qualitative results in Fig. 4 further illustrate this point. In Fig. 4-(a), a scenario where patches are randomly selected and directional loss based on the text direction is applied without incorporating all proposed losses and modules, demonstrates a significant limitation. Here, the entire source image undergoes stylization, which leads to a distortion of the semantic content originally present in the image. Fig. 4-(b) demonstrates that applying only $\mathcal{L}_{\text{dir}}$ leads to focused stylization on the targeted object, enabled by TMPS, but also results in notable background changes and some loss or alteration of object details. In contrast, Fig. 4-(c), where $\mathcal{L}_{\text{dir}}$ is combined with $\mathcal{L}_{\text{con}}$, effectively preserves crucial object details like the chair’s shadow and shape of the hat, though background alterations remain. The comparison of Fig. 4-(d) and (e) reveals the impact of the $\mathcal{L}_{\text{con}}$; in (d), despite achieving object-centric stylization, there is a noticeable loss in object details, particularly in cropped areas. Meanwhile, the comparison of Fig. 4-(c) and (e) demonstrates that with the full complement of losses, both the background remains unchanged and the object details are preserved without distortion, highlighting the effectiveness of the ABP loss.
4.5 Complex input texts
We have conducted further experiments using a diverse set of examples featuring intricate ‘source’ and ‘style’ texts results in Fig. 5. For example, in the ‘cake to emerald’ scenario, although the stylized image with simple text retains aspects of the original cake’s style, the detailed source text enables the creation of a cake styled purely in emerald. Similarly, in the ‘barn to snowy’ scenario, the model adeptly preserves the background’s style while effectively applying stylization to the foreground object. Furthermore, our experiments incorporating complex ‘style’ texts illustrate the complexity of the text does not impede our method’s ability to achieve the intended stylization effects. These results clearly demonstrate that our method is adept not only at managing simple texts but is equally effective with complex textual inputs. This expansion of our testing framework underscores our method’s robustness and versatility in accommodating a broad spectrum of textual complexities.
@@ 5 Conclusion
In this paper, we have presented Text-driven Object-Centric Style Transfer (TEXTOC), a new approach for transforming the appearance of objects in images based purely on textual descriptions. This method eliminates the need for reference style images or segmentation masks, facilitating complex and tailored stylizations directly from text. At the heart of TEXTOC are three pivotal components: the Text-Matched Patch Selection with Pre-fixed Region Selection module, the Patch-wise Co-Directional (PCD) loss and the Adaptive Background Preservation (ABP) loss. The TMPS and PRS modules are pivotal in accurately identifying the location of objects linked to the source text. The PCD loss, complemented by our Text-Matched Patch Selection (TMPS), precisely identifies and selects patches for stylization, ensuring an artifact-free style transfer through consistent CLIP-embedding distribution. Meanwhile, the ABP loss plays a critical role in preserving the original style and structure of the background. It adeptly adjusts to dynamically identified background areas, effectively preventing unintended alterations during the style transfer process. Extensive experiments validate the superior performance of our TEXTOC, demonstrating its capability to achieve high-quality, text-driven object-centric stylization while preserving the overall visual coherence and integrity of images.
@@
Appendix A Detailed training process of the TEXTOC
@@
  • The model takes three inputs: a source image ($I^{src}$), which is the subject of the style transfer; a source text ($T^{src}$), which identifies the specific object in the source image to be modified; and a style text ($T^{sty}$), describing the desired style to be applied to the object.
@@ A.1 Pre-fixed Region Selection (PRS)
In our approach, we process the source image ($I^{src}$) by dividing it into a $9\times 9$ grid. From each grid point, we generate three patches of varying sizes, centered on these points, resulting in a total of 243 patches. These patches are then passed through the Text-Matched Patch Selection (TMPS) module, which identifies and selects the patches that are most relevant to the source text ($T^{src}$). We use a specific threshold to determine the foreground region ($M^{\text{fg}}$) of the image: grids that exceed this threshold in terms of patch relevance are classified as part of the foreground. After the initial iterations, further patch generation is concentrated within this coarse foreground region. The number of patches generated in subsequent iterations is adjusted proportionally to the count of grids identified as part of the foreground region. This method ensures a focused and relevant application of style transfer where it is most pertinent according to the text input.
Appendix B Comprehensive Analysis of our TEXTOC
In our study, we conduct experiments specifically designed to evaluate the impact of the patch distribution consistency loss within our TEXTOC model. The primary objective of these experiments was to ascertain how this loss function influences the overall effectiveness of the model in achieving high-quality stylization. Additionally, we present the stylization results at various stages of the model’s training process. This presentation is instrumental in demonstrating the progressive impact of the proposed loss function. Through these results, we aim to showcase how the patch distribution consistency loss enhances the model’s capability to deliver consistently high-quality stylizations. This detailed examination and presentation of results not only highlight the efficacy of the patch distribution consistency loss in our TEXTOC model but also provide valuable insights into the model’s operational dynamics throughout its training phase.
@@ B.1 The patch distribution consistency loss
To demonstrate the impact of the patch distribution consistency loss, denoted as $\mathcal{L}_{\text{con}}$, we conduct an ablation study. The findings from this study are displayed in Fig. 8. A key observation from our experiments is that the inclusion of $\mathcal{L}_{\text{con}}$ significantly contributes to the preservation of vital features in the images, even after the style transformation process. Notable examples include the retention of text on a T-shirt, the intricate pattern of a tropical fish, and the clarity of numbers on a bowling ball. These results highlight that the implementation of the $\mathcal{L}_{\text{con}}$ loss is crucial in maintaining a focused distribution of patches on the object. This focused approach is what enables the preservation of essential details and textures during the style transformation, ensuring that the core visual elements of the object remain intact and recognizable post-transformation. This study thus underscores the effectiveness of the $\mathcal{L}_{\text{con}}$ loss in enhancing the quality and fidelity of object-centric style transfers in our TEXTOC model.
Figure 8: Qualitative comparison demonstrating the effect of the PCD loss by contrasting results with and without the $\mathcal{L}_{\text{con}}$ loss.
B.2 Visualization of style transfer process
Fig. 9 shows the iterative process of style transfer as conducted by our model. Initially, the model prioritizes the preservation of the background, ensuring it remains as close to the original image as possible. This early focus on background fidelity is a crucial step in maintaining the overall integrity and context of the source image. As the training progresses through successive iterations, the model begins to refine its approach. It gradually learns to enhance and accentuate the details within both the object and the background. This progression illustrates the model’s sophisticated capability to strike a balance between maintaining background fidelity and enhancing the object of interest. This iterative learning and adaptation process is a key aspect of our model’s functionality. It demonstrates how the model evolves to effectively manage the complexities of style transfer, ensuring that both the object’s details and the background’s essence are harmoniously preserved and enhanced.
Figure 9: Visualization of the style transfer process at intervals of every 10 iterations. This sequence illustrates the progressive transformation and refinement of the image style over time. The parts marked with blue dashed lines denote the iterations where the Pre-fixed Region Selection module operates, and the green patches in each iteration denote the patches selected through the Text-Matched Patch Selection. Areas not included in the patch are the regions where the Adaptive Background Preservation loss ($\mathcal{L}_{\text{abp}}$) is applied.
@@ B.3 Computational overhead of TMPS and PRS
Appendix C Limitations of TEXTOC

### v3

@@ 4.1 Implementation Details
We commence our training with an initial set of 20 early iterations and completed a total of 200 iterations using the Adam optimizer [32]. The training began with an initial learning rate of $5\times 10^{-4}$, halved after the first 100 iterations. We incorporate the perspective augmentation function from the PyTorch library [44]. Notably, our method processes each source image independently, using only the source image, source text, and style text, without requiring additional training data. The average training time for each text prompt is approximately 45 seconds on an NVIDIA A6000 GPU. For further details on our training procedure, please refer to Appendix A.
@@
4.2 Evaluation metrics
[…]
$\begin{gathered}\text{Sim}_{F}=\frac{E_{T}\left(T^{\text{tgt}}_{i}\right)\cdot E_{I}(I^{\text{out}}\odot M^{\text{fg\_gt}})}{|E_{T}\left(T^{\text{tgt}}_{i}\right)|\cdot|E_{I}(I^{\text{out}}\odot M^{\text{fg\_gt}})|},\\ \text{L1}_{B}=\frac{\sum|I^{\text{src}}\odot M^{\text{bg\_gt}}-I^{\text{out}}\odot M^{\text{bg\_gt}}|}{255\cdot\text{num}(I^{\text{src}}\odot M^{\text{bg\_gt}})},\end{gathered}$ (7)
@@ 4.3 Comparison to state-of-the-arts
In the qualitative evaluation, the first row of Fig. 4 showcases the style editing of a mountain into a volcano using the style text “Volcano”. Our method adeptly preserves the inherent structure of the mountain while seamlessly integrating the volcanic style. In contrast, Text2LIVE often reconstructs an entirely new mountain structure, typically characterized by a singular hole indicative of magma. This comparison underscores the nuanced capability of our approach in maintaining the original form while applying distinct style editing. Utilizing source texts like “wooden door”, “croissant” and “red umbrella”, our method demonstrates a precise capture of object locations, contrasting with other models that often indiscriminately apply style editing to the entire image, leading to unfavorable background metric scores. Additionally, the six diffusion-based style editing models frequently alter the target object, resulting in poor $Con_{F}$ scores. While Text2LIVE shows results similar to ours, it notably applies style editing to unintended areas (e.g., wooden door with “gray marble” as style text) and fails to properly reflect the style indicated by the style text (e.g., red umbrella in ”Starry night by Vincent Van Gogh”).
@@ 4.4 Ablation study
[…]
The qualitative results in Fig. 5 further illustrate this point. In Fig. 5-(a), a scenario where patches are randomly selected and directional loss based on the text direction is applied without incorporating all proposed losses and modules, demonstrates a significant limitation. Here, the entire source image undergoes style editing, which leads to a distortion of the semantic content originally present in the image. Fig. 5-(b) demonstrates that applying only $\mathcal{L}_{\text{dir}}$ leads to focused style editing on the targeted object, enabled by TMPS, but also results in notable background changes and some loss or alteration of object details. In contrast, Fig. 5-(c), where $\mathcal{L}_{\text{dir}}$ is combined with $\mathcal{L}_{\text{con}}$, effectively preserves crucial object details like the chair’s shadow and shape of the hat, though background alterations remain. The comparison of Fig. 5-(d) and (e) reveals the impact of the $\mathcal{L}_{\text{con}}$; in (d), despite achieving object-centric style editing, there is a noticeable loss in object details, particularly in cropped areas. Meanwhile, the comparison of Fig. 5-(c) and (e) demonstrates that with the full complement of losses, both the background remains unchanged and the object details are preserved without distortion, highlighting the effectiveness of the ABP loss.
4.5 Comparison to mask-guided generative models
To highlight the differences from mask-based models and showcase the natural style-editing capabilities of our Style-Editor model, we conduct a comparative analysis with existing mask-guided generative models [1, 41]. These models typically rely on masks to specific areas for image editing. Our comparison, as depicted in Fig. 6, highlights a key distinction. Unlike mask-based models, which may inadvertently distort vital details of the object, Style-Editor consistently maintains the structural integrity of the object, focusing solely on altering its style. This process is crucial in preserving the overall coherence and authenticity of the image. Moreover, Style-Editor effectively identify relevant objects based on the provided text and apply style edits directly to these areas, eliminating the need for labor-intensive and sometimes imprecise manual mask creation. Style-Editor, thus, not only preserves the integrity of the objects but also enhances efficiency, offering a more user-friendly alternative to traditional mask-guided generative models.
Figure 6: Comparative analysis of Style-Editor with other generative models using mask input. This figure demonstrates the superiority of our model in producing enhanced results compared to methods that use object masks as guidance for style editing tasks.
@@ 5 Conclusion
In this paper, we have presented Text-driven Object-Centric Style editing (Style-Editor), a new approach for editing the appearance of objects in images based purely on textual descriptions. This method eliminates the need for reference style images or segmentation masks, facilitating complex and tailored style editing directly from text. At the heart of Style-Editor are three pivotal components: the Text-Matched Patch Selection with Pre-fixed Region Selection module, the Patch-wise Co-Directional (PCD) loss and the Adaptive Background Preservation (ABP) loss. The TMPS and PRS modules are pivotal in accurately identifying the location of objects linked to the source text. The PCD loss, complemented by our Text-Matched Patch Selection (TMPS), precisely identifies and selects patches for style editing, ensuring an artifact-free style editing through consistent CLIP-embedding distribution. Meanwhile, the ABP loss plays a critical role in preserving the original style and structure of the background. It adeptly adjusts to dynamically identified background areas, effectively preventing unintended alterations during the style editing process. Extensive experiments validate the superior performance of our Style-Editor, demonstrating its capability to achieve high-quality, text-driven object-centric style editing while preserving the overall visual coherence and integrity of images.
Acknowledgments
This work was supported by the 2025 innovation base artificial intelligence data convergence project project with the funding of the 2025 government (Ministry of Science and ICT) (S2201-24-1002). The supercomputing resources for this work was supported by Grand Challenging Project of Supercomputing AI Education and Research Center, DGIST.
@@ Acknowledgments
Appendix A Detailed training process of the Style-Editor
@@
  • The model takes three inputs: a source image ($I^{src}$), which is the subject of the style editing; a source text ($T^{src}$), which identifies the specific object in the source image to be modified; and a style text ($T^{sty}$), describing the desired style to be applied to the object.
@@ A.1 Pre-fixed Region Selection (PRS)
In our approach, we process the source image ($I^{src}$) by dividing it into a $9\times 9$ grid. From each grid point, we generate three patches of varying sizes, centered on these points, resulting in a total of 243 patches. These patches are then passed through the Text-Matched Patch Selection (TMPS) module, which identifies and selects the patches that are most relevant to the source text ($T^{src}$). We use a specific threshold $\tau$ (we set as 2) to determine the foreground region ($M^{\text{fg}}$) of the image: grids that exceed this threshold in terms of patch relevance are classified as part of the foreground. After the initial iterations, further patch generation is concentrated within this coarse foreground region. The number of patches generated in subsequent iterations is adjusted proportionally to the count of grids identified as part of the foreground region. This method ensures a focused and relevant application of style editing where it is most pertinent according to the text input.
Appendix B Comprehensive Analysis of our Style-Editor
In this section, we provide a detailed analysis of our Style-Editor model by evaluating its components and performance through various experiments. Specifically, we focus on the impact of the patch distribution consistency loss, visualize the style editing process, assess the computational overhead of the TMPS and PRS modules, explore the effect of patch size, and compare our patch-wise approach with segmentation-based methods. This detailed examination and presentation of results not only highlight the efficacy of our Style-Editor model but also provide valuable insights into the model’s operational dynamics throughout its training phase.
@@ B.1 The patch distribution consistency loss
To demonstrate the impact of the patch distribution consistency loss, denoted as $\mathcal{L}_{\text{con}}$, we conduct an ablation study. The findings from this study are displayed in Fig. 9. A key observation from our experiments is that the inclusion of $\mathcal{L}_{\text{con}}$ significantly contributes to the preservation of vital features in the images, even after the style editing process. Notable examples include the retention of text on a T-shirt, the intricate pattern of a tropical fish, and the clarity of numbers on a bowling ball. These results highlight that the implementation of the $\mathcal{L}_{\text{con}}$ loss is crucial in maintaining a focused distribution of patches on the object. This focused approach is what enables the preservation of essential details and textures during the style editing, ensuring that the core visual elements of the object remain intact and recognizable post-transformation. This study thus underscores the effectiveness of the $\mathcal{L}_{\text{con}}$ loss in enhancing the quality and fidelity of object-centric style editings in our Style-Editor model.
Figure 9: Qualitative comparison demonstrating the effect of the $\mathcal{L}_{\text{pcd}}$ loss by contrasting results with and without the $\mathcal{L}_{\text{con}}$ loss.
B.2 Visualization of style editing process
Fig. 10 shows the iterative process of style editing as conducted by our model. Initially, the model prioritizes the preservation of the background, ensuring it remains as close to the original image as possible. This early focus on background fidelity is a crucial step in maintaining the overall integrity and context of the source image. As the training progresses through successive iterations, the model begins to refine its approach. It gradually learns to enhance and accentuate the details within both the object and the background. This progression illustrates the model’s sophisticated capability to strike a balance between maintaining background fidelity and enhancing the object of interest. This iterative learning and adaptation process is a key aspect of our model’s functionality. It demonstrates how the model evolves to effectively manage the complexities of style editing, ensuring that both the object’s details and the background’s essence are harmoniously preserved and enhanced.
Figure 10: Visualization of the style editing process at intervals of every 10 iterations. This sequence illustrates the progressive transformation and refinement of the image style over time. The parts marked with blue dashed lines denote the iterations where the Pre-fixed Region Selection (PRS) module operates, and the green patches in each iteration denote the patches selected through the Text-Matched Patch Selection (TMPS). Areas not included in the patch are the regions where the Adaptive Background Preservation loss ($\mathcal{L}_{\text{abp}}$) is applied.
@@ B.3 Computational overhead of TMPS and PRS
B.4 Impact of Patch Size
We conducted additional experiments to investigate the impact of patch size in the TMPS module. The experimental setup and results are presented in Tab. 4 and Tab. 5, highlighting how variations in patch sizes affect key performance metrics. Our findings demonstrate that the TMPS module remains robustness across different patch sizes. Notably, our adaptive selection of patch sizes, ranging from 64 to 128 for images of 512$\times$512 size, yields well-rounded results throughout our experiments. This adaptive approach reflects a careful trade-off: smaller patches enhance image quality metrics, such as PSNR, $\text{Con}_{F}$, and $\text{Con}_{B}$, while larger patches help preserve stylization consistency, especially around object boundaries. Nevertheless, it is worth noting that such a reduction in patch size might also decrease the CLIP similarity for the foreground ($\text{Sim}_{F}$), potentially compromising the content’s integrity.
Foreground quality metrics Background quality metrics
[…]
patch 64 0.3 3.46 0.09 1.2 0.11 0.90 0.08 27.92
patch 96 0.32 3.94 0.09 1.25 0.12 0.89 0.08 27.56
patch 128 0.33 4.24 0.10 1.29 0.13 0.89 0.08 27.17
Ours (patch 64-128) 0.33 3.75 0.10 1.15 0.10 0.90 0.07 27.65
Table 4: Quantitative evaluation of fixed patch sizes. Metrics include similarity (SimF, L1B), content loss (ConF, ConB), style loss (StyB), structural similarity index (SSIM), perceptual distance (DISTS), and peak signal-to-noise ratio (PSNR). $\uparrow$ indicates higher values are better, while $\downarrow$ indicates lower values are better.
Foreground quality metrics Background quality metrics
[…]
patch 32-64 0.29 2.90 0.09 1.12 0.11 0.91 0.08 28.41
Ours (patch 64-128) 0.33 3.75 0.10 1.15 0.10 0.90 0.07 27.65
patch 128-256 0.34 4.43 0.11 1.37 0.15 0.87 0.09 26.05
Table 5: Quantitative evaluation of adaptive patch sizes. Metrics include similarity (SimF, L1B), content loss (ConF, ConB), style loss (StyB), structural similarity index (SSIM), perceptual distance (DISTS), and peak signal-to-noise ratio (PSNR). $\uparrow$ indicates higher values are better, while $\downarrow$ indicates lower values are better.
B.5 Comparison with Segmentation method
To evaluate the effectiveness of our approach compared to segmentation-based methods, we conducted experiments by applying a mask to our results using the open-vocabulary segmentation model ODISE [61]. The comparative results are illustrated in Fig. 11. When applying a ‘Sunlight’ style to a towel image, our model generates results where light naturally diffuses around the object. In contrast, segmentation-based methods rely on per-pixel binary classification (0 or 1) to determine the presence of an object, resulting in abrupt and unnatural transitions at the boundaries where the style is applied. This disrupts the overall continuity and realism of the stylized image. Furthermore, segmentation-based methods typically require additional training on datasets, such as the MSCOCO [38] dataset. These findings highlight the advantage of our model in maintaining seamless and natural style editing, avoiding the artifacts commonly observed at object boundaries in segmentation-based methods. Moreover, by leveraging the CLIP encoder for both style transfer and object patch identification, our approach eliminates the need for pre-trained segmentation networks, achieving both simplicity and effectiveness.
Figure 11: Qualitative comparison between our results and our results with segmentation masks applied.
Appendix C Visualization of evaluation metric
To evaluate the performance of object-centric style editing, we conducted evaluations using the annotations from the MS COCO 2017 dataset. Fig. 12 presents the visualized results used as reference examples for the evaluation metrics employed in our experiments. In particular, $M^{\text{fg\_gt}}$ denotes the ground truth (GT) mask corresponding to the target class. Using this mask, we calculated the foreground quality metrics by masking regions outside the class object areas and cropping the images according to the mask area, as shown in $I^{\text{src}}\odot M^{\text{fg\_gt}}$ and $I^{\text{out}}\odot M^{\text{fg\_gt}}$ in the Fig. 12. Conversely, the background quality metrics were measured using images where the object regions were excluded, as represented by $I^{\text{src}}\odot M^{\text{bg\_gt}}$ and $I^{\text{out}}\odot M^{\text{bg\_gt}}$ in the Fig. 12.
Figure 12: Example images for evaluating foreground quality metrics and background quality metrics.
Appendix D Complex input texts
We have conducted further experiments using a diverse set of examples featuring intricate ‘source’ and ‘style’ texts resulting in Fig. 13. For example, in the ‘cake to emerald’ scenario, although the stylized image with simple text retains aspects of the original cake’s style, the detailed source text enables the creation of a cake styled purely in emerald. Similarly, in the ‘barn to snowy’ scenario, the model adeptly preserves the background’s style while effectively applying style editing to the foreground object. Furthermore, our experiments incorporating complex ‘style’ texts illustrate the complexity of the text does not impede our method’s ability to achieve the intended editing effects. These results clearly demonstrate that our method is adept not only at managing simple texts but is equally effective with complex textual inputs. This expansion of our testing framework underscores our method’s robustness and versatility in accommodating a broad spectrum of textual complexities.
Figure 13: Our style editing results with complex image-text scenarios.
Appendix E Limitations of Style-Editor